# IDEA: An Inverse Domain Expert Adaptation Based Active DNN IP Protection Method

### v1

@@ V-B4 Verification and tracing performance
  Method Dataset $\rightarrow$ CIFAR-10 CIFAR-100 Caltech-101 GTSRB ImageNet-10
Setting $\downarrow$ authorized benign authorized benign authorized benign authorized benign authorized benign
  DeepIPR [19] - 93.07 (-0.51) 7.82 72.08 (-1.23) 2.56 74.89 (-0.84) 4.30 97.55 (+0.26) 11.75 87.82 (-1.03) 28.51
SSAT [21] $y^{\prime}=y+1$ 93.53 (-0.05) 0.60 72.95 (-0.36) 0.21 76.04 (+0.31) 0.81 97.85 (+0.56) 0.41 87.57 (-1.28) 2.10
DSN [24] - 93.06 (-0.52) 11.55 72.42 (-0.89) 2.22 77.11 (+1.38) 1.10 97.97 (+0.68) 1.51 89.04 (+0.19) 10.96
Pixel Shuffling [18] $b$ = 2 92.41 (-1.17) 44.71 67.53 (-5.78) 10.87 73.48 (-2.25) 41.27 96.61 (-0.68) 8.52 86.89 (-1.96) 58.45
$b$ = 4 89.28 (-4.30) 23.78 57.48 (-15.83) 3.45 64.84 (-10.89) 17.46 94.46 (-2.83) 8.80 84.93 (-3.92) 56.50
$b$ = 8 82.36 (-11.22) 20.07 42.27 (-31.04) 5.33 53.14 (-22.59) 5.24 88.19 (-9.10) 5.09 83.37 (-5.48) 48.93
@@
Ours $\mathcal{L}_{\text{kl}}$, $\mathcal{L}_{\text{at}}$, $\mathcal{L}_{\text{crd}}$ 93.15 (-0.43) 10.13 73.10 (-0.21) 1.33 77.23 (+1.50) 1.04 97.78 (+0.49) 2.01 89.31 (+0.46) 10.35
@@ V-C Active Protection Methods Comparison
We re-implemented four recent active protection methods, DeepIPR [19], SSAT [21], DSN [24], and pixel shuffling [18], on ResNet-18 across the five datasets for comparison. We also directly distilled a student model from the MoE model with only KL-divergence and without attention and contrastive representation loss terms. We call this model the “simple-distilled model”. Table VI presents the accuracies of the protected models on authorized and benign test images. The accuracy change for each protected model and dataset is also provided in bracket.
The performance of the pixel shuffling [18] method is dependent on the shuffling block size $b$, which trades model’s accuracy on unauthorized images for model’s accuracy on authorized images. For low-resolution datasets, $b=2$ provides a good trade-off between effectiveness and fidelity. For ImageNet-10, it can only keep the prediction accuracy of the unauthorized images to below 50% by degrading the accuracy of authorized images by more than 5% with optimal $b=8$.
DeepIPR [19], SSAT [21], and DSN [24] can keep the accuracy drop to lower than 1.5% on authorized images and sufficiently degrade the performance on unauthorized images (<40%). However, these three methods have their own security vulnerability. The passport layers of DeepIPR can be easily localized since they require additional inputs. The attacker can replace these passport layers with custom normalization layers and re-train them to achieve high accuracy on benign images without passport inputs, as demonstrated in [51]. SSAT assigns ground truth labels to poisoned training samples and wrong labels to clean training samples. The ground-truth label $y$ of clean input image is mapped to the wrong label $y^{\prime}=y+1$. This fixed bijective mapping can be easily reverse-engineered by observing the model’s outputs to a small set of labeled test images. Any other deterministic bijective mapping is as vulnerable. The DSN method directly pastes a fixed binary pattern on the test images as secret key, which can be easily spotted by human eyes.
IDEA achieves good effectiveness and fidelity on all datasets. Layer substitution attack is not able to break IDEA as we make no modification to the model architecture. Query-based reverse-engineering is also infeasible as the outputs of unauthorized images are randomized. Since the perturbations introduced by SteganoGAN encoder are imperceivable and sample-specific, it is impossible for the attacker to recognize and copy the secret key.
Although the simple-distilled model can achieve similar classification accuracy as the protected model, their differences will be more closely scrutinized in Sec. VI.
@@ VI A Closer Look at Latent Representations
In this section, the hidden layer outputs of the unprotected, MoE, simple-distilled and protected models are examined from three perspectives: MI estimation, attention map visualization, and representation visualization.
@@ VI-A MI Estimation
Fig. 6 visualizes the estimated MI between the latent representations collected from paired source (authorized) and target (unauthorized) test samples, with the first row showing the MI between authorized and benign images, and the second row showing the MI between authorized and noise images. Layers 1 to 4 are the selected layers for MI minimization and multi-layer attention and contrastive representation knowledge distillation, progressing from the shallow to deep layer.
For all selected layers, we observe that the estimated MI on the unprotected model is consistently high. This is expected, as this model has not been trained to recognize steganographic perturbations and thus produces similar representations for all types of images. In contrast, the MI between the source and target images estimated on the MoE model is significantly lower than that on the unprotected model, demonstrating that the training objectives in Eqs. (18) and (19) effectively increase the expected risk of the two fake experts. The protected model also shows similarly low MI values, indicating that the protected model faithfully emulates the MoE model’s behavior across all selected layers. Consequently, if the input images are not embedded with the correct key, the adversary can only obtain meaningless intermediate layer representations that are comparable to those output by the fake experts. However, the MI estimated on the simple-distilled model is higher than that of the MoE and protected models, and in some cases even higher than that of the unprotected model. This indicates that using only KL-divergence constraints on the output layer during knowledge distillation is insufficient to protect the model functionalities. Specifically, the shallow layers of the simple-distilled model exhibit good generalizability on the target domains, making it easier to break the active authorization protection through fine-tuning, as these well-performing shallow layers require almost no adjustment.
Figure 7: Visualization of attention maps generated by the original unprotected, MoE, protected and simple-distilled models, respectively. The test images are randomly selected from ImageNet-10.
Figure 8: t-SNE visualization of the latent representations of the original unprotected (top left), simple-distilled (bottom left), MoE (top right) and protected (bottom right) models obtained from 1500 ImageNet-10 test images. The latent representations of the benign/authorized/noise test images are shown as blue squares, green dots, and red triangles, respectively.
@@ VI-B Attention Map Visualization
Fig. 7 shows the attention maps generated at the output of selected layers by feeding paired benign/authorized/noise test images into the four models. It shows that the unprotected model outputs similar attention maps for all three domains of test images at all selected layers. Conversely, the MoE model produces similar attention maps to the unprotected model only on input of authorized images. On input of benign or noise images, the MoE model exhibits attention behaviors akin to an untrained model to these inputs. Moreover, the protected model perfectly imitates the intermediate layer behaviors of the MoE model, as it outputs almost identical attention maps as those of the MoE model for the same domains of test images at all the selected layers. However, the simple-distilled model exhibits similar attention maps for all types of images at the first three selected layers, and these attention maps are similar to those of the unprotected model at the same layers.
@@
We randomly selected 500 benign images from the ImageNet-10 dataset and generated their authorized and noise versions to create a mixed test set with 1500 images. These test images were separately fed into the four models, and their outputs at the four selected layers were collected. The latent representations of the three types of test images are visualized using the t-SNE [38] in Fig. 8. For the unprotected model, the points corresponding to the paired benign/authorized/noise images almost completely overlap across all selected layers. This further illustrates that the source and target images have high estimated MI at the hidden layers. The simple-distilled model has similar behaviors in layers 1 to 3, but the representations in layer 4 form three distinct clusters. This result, again, reveals that the simple-distilled model has good generalizability at the shallow layers, and the representations of the three types of images are indistinguishable until the deep layers. On the contrary, for the MoE and protected models, the points corresponding to benign/authorized/noise images are well clustered across all selected layers. This indicates that the MI minimization has successfully made the two fake experts behave differently from the real expert, and the protected model effectively mimics the behavior of the MoE model at not only the output layer but also the selected hidden layers.
@@ VII Security Analysis
We further investigate the robustness of IDEA against both model transformation and reverse-engineering attacks. Model transformation attacks aim to remove the active authorization protection by directly modifying the pirated model through fine-tuning, model-pruning, and transfer-learning. Reverse-engineering attacks aim to reverse engineer the encoding process without modifying the pirated model.
Model Transformation Attacks
TABLE VII: Accuracy (%) of fine-tuned models on benign test datasets. The last column lists the baseline classification accuracy of the unprotected models on benign test datasets.
  Dataset Ratio Protected model Simple-distilled model baseline
FTAL FTLL RTAL RTLL FTAL FTLL RTAL RTLL
  CIFAR-10 10 % 58.38 57.00 60.58 58.46 82.27 83.62 87.26 69.51 93.58
20 % 60.82 58.98 62.01 65.74 88.52 85.47 91.60 73.58
30 % 61.54 58.73 64.86 66.18 90.23 86.33 92.28 75.92
CIFAR-100 10 % 45.73 41.91 26.46 22.39 58.36 41.92 22.32 20.28 93.31
20 % 47.12 41.54 40.83 32.51 66.02 41.59 41.67 32.00
30 % 49.00 42.46 50.12 39.47 68.08 41.69 47.38 36.53
Caltech-101 10 % 53.95 52.80 38.16 32.91 60.00 51.47 33.72 31.99 75.73
20 % 55.27 52.05 48.47 39.31 62.88 53.31 39.02 35.62
30 % 56.48 52.80 57.69 42.54 63.63 55.62 48.36 40.35
GTSRB 10 % 50.78 47.28 11.08 9.33 68.31 53.82 12.10 9.64 97.29
20 % 54.13 46.17 20.51 14.49 89.52 54.08 20.85 16.20
30 % 56.37 41.52 29.95 20.07 92.62 48.51 30.48 22.22
ImageNet-10 10 % 64.35 70.19 60.77 62.69 78.27 80.08 85.38 78.08 88.85
20 % 72.77 70.46 70.81 71.00 81.65 81.88 88.00 79.31
30 % 72.35 70.54 69.88 73.04 83.00 83.04 88.96 80.12
Fine-tuning. Following the configurations in [52], we apply the following four fine-tuning strategies, fine-tune all layers (FTAL), fine-tune last layer (FTLL), re-train all layers (RTAL), and retrain last layer (RALL) on the protected model.
  • FTAL: The weights in all the layers are updated.
  • FTLL: Only the weights in the last layer are updated.
  • RTAL: The weights in the last layer are randomly initialized, and then the weights in all the layers are updated.
  • RTLL: The weights in the last layer are randomly initialized, but only the weights in the last layer are updated.
The protected model is fine-tuned for 30 epochs using 10%, 20%, and 30% of the samples from the benign training dataset. For comparison, we also conduct the same attacks on the simple-distilled model.
Table VII shows the accuracy of the fine-tuned models on the benign test dataset. The baseline performance of the unprotected models on the benign test dataset is shown in the last column. For fine-tuned protected models, the performance improves with increasing ratio of accessible samples. However, even with 30% training samples, the inference accuracy on benign data is still significantly lower than the baseline performance. In most cases, the model fine-tuned from the simple-distilled model has higher accuracy than from the protected model, especially for FTAL. This result agrees with our observation in Sec. VI that the shallow layers of the simple-distilled model can be exploited for fine-tuning attack. Therefore, the multi-layer attention and contrastive representation losses are required to prevent the distilled shallow layers from leaking the real expert functionality of the teacher model.
@@
Model-pruning. Weight pruning (WP) and filter pruning (FP), are commonly used to reduce model size by removing parameters that contribute insignificantly to inference performance [53]. WP globally prunes a fraction (ranging from 5% to 95%, with a step size of 5%) of the smallest magnitude weights. FP iteratively prunes filters in a specific convolutional layer, starting from the smallest magnitude to the largest. Following the configuration in [54], we select the last convolutional layer of ResNet-18 as the target layer for FP implementation.
@@
TABLE VIII: Accuracy (%) of models trained using transfer-learning from unprotected and protected models. The original training datasets are listed in the first row and the private dataset to be adapted in the first column. ImageNet-10$\dagger$ is another subset of ImageNet consisting of 10 classes that do not overlap with ImageNet-10.
  Dataset CIFAR-10 CIFAR-100 Caltech-101 GTSRB ImageNet-10$\dagger$
unprotected protected unprotected protected unprotected protected unprotected protected unprotected protected
  CIFAR-10 - - 30.41 17.39 37.35 26.05 50.85 22.80 - -
CIFAR-100 75.25 61.47 - - 55.50 30.14 65.69 36.66 - -
Caltech-101 68.17 54.62 41.73 26.13 - - 63.06 43.95 - -
GTSRB 46.78 35.43 18.70 10.17 31.99 20.00 - - - -
ImageNet-10 - - - - - - - - 50.34 40.36
Transfer-learning. This attack aims to utilize the well-trained feature extractor to adapt the pirated model to a different task using a private dataset. The common transfer-learning paradigm freezes all layers except the last fully connected layer, which is randomly initialized based on the new task. For the four low-resolution datasets, we perform transfer learning on both unprotected and protected models using other low-resolution datasets. For the high-resolution ImageNet-10, we select another 10 classes from the entire ImageNet dataset to form a non-overlapping subset ImageNet-10$\dagger$ for transfer-learning. The results in Tabel VIII show that models that are transfer-learnt from protected models exhibit worse performance than from unprotected models as protected models output meaningless features by imitating fake experts.
TABLE IX: Accuracy (%) of pirated models on reverse-engineered images generated based on Assumptions 1 and 2.
@@
Reverse-engineering Attacks
@@
Assumption 1: The attacker only has a small portion of the benign images to train a generator $\mathcal{G}(\cdot)$ to produce encoded images for the pirated model. The optimization objective of the generator can be expressed as follows:
$\mathop{\arg\min}\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{sub}}}\left[\mathcal{L}_{\text{CE}}(f_{\theta^{p}_{j}}(\mathcal{G}(x)),y)\right],$ (29)
where $D_{\text{sub}}$ denotes the benign subset hold by the attacker.
@@
Assumption 2: The attacker has not only a small portion of benign images but also their authorized versions embedded with the correct key by eavesdropping on legitimate users. This scenario enables the attacker to train the generator more efficiently by simultaneously minimizing the MSE loss between the generated and authorized images as follows:
$\mathop{\arg\min}\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{sub}}}\Big{[}\mathcal{L}_{\text{CE}}(f_{\theta^{p}_{j}}(\mathcal{G}(x)),y)+\lambda_{4}\mathcal{L}_{\text{MSE}}(\mathcal{G}(x),x^{+}_{j})\Big{]},$ (30)
where $\lambda_{4}$ is a weighting factor.
@@
We simulate these attacks by training a generator with different numbers of benign images or benign-authorized pairs, ranging from 500 to 5000. The generator has a similar architecture to the SteganoGAN encoder but without the key input port. Both its input and output are 3-channel color images of the same size. Table IX reports the accuracy of the pirated model on test images created by the generator. Under Assumption 1, most reverse-engineering attacks result in low prediction accuracy close to random guessing. The best attack performance achieved on Caltech-101 with 5000 available benign images is merely 22.13%. As expected, the attack performance under Assumption 2 is much better in most cases, with an average improvement of 9.55%. However, even with 5000 available benign-authorized pairs, which are difficult for the attacker to collect in reality, the best attack performance of 56.84% achievable on CIFAR-10 is still far lower than the baseline. Therefore, IDEA is resilient against even the strongest assumption of reverse-engineering attacks.
VIII Conclusion

### v2

@@ V-B4 Licensee Verification
  Method Dataset $\rightarrow$ CIFAR-10 Caltech-101 GTSRB ImageNet-10
Setting $\downarrow$ Authorized Benign Authorized Benign Authorized Benign Authorized Benign
  DeepIPR [19] - 93.07 (-0.51) 7.82 74.89 (-0.84) 4.30 97.55 (+0.26) 11.75 88.82 (-0.03) 38.96
SSAT [21] $y^{\prime}=y+1$ 93.53 (-0.05) 0.60 76.04 (+0.31) 0.81 97.85 (+0.56) 0.41 87.57 (-1.28) 2.10
DSN [26] - 93.06 (-0.52) 11.55 77.11 (+1.38) 1.10 97.97 (+0.68) 1.51 89.04 (+0.19) 10.96
Pixel Shuffling [18] $b=2$ 92.41 (-1.17) 44.71 73.48 (-2.25) 41.27 96.61 (-0.68) 8.52 86.89 (-1.96) 58.45
$b=4$ 89.28 (-4.30) 23.78 64.84 (-10.89) 17.46 94.46 (-2.83) 8.80 84.93 (-3.92) 56.50
$b=8$ 82.36 (-11.22) 20.07 53.14 (-22.59) 5.24 88.19 (-9.10) 5.09 83.37 (-5.48) 48.93
@@
Ours $\mathcal{L}_{\text{kl}}+\mathcal{L}_{\text{at}}+\mathcal{L}_{\text{crd}}$ 93.15 (-0.43) 10.13 77.23 (+1.50) 1.04 97.78 (+0.49) 2.01 89.31 (+0.46) 10.35
@@ V-C Active Protection Methods Comparison
Four recent active protection methods, DeepIPR [19], SSAT [21], DSN [26], and pixel shuffling [18], are re-implemented on ResNet-18 across the five datasets for comparison. We also directly distilled a student model from the MoE model with only KL-divergence and without attention and contrastive representation loss terms. We call this model the “elementary-distilled” model. Table V presents the accuracies of the protected models on authorized and benign test images. The difference in accuracy between the authorized and benign test images for each protected model and dataset is also provided in bracket.
The performance of the pixel shuffling [18] method is dependent on the shuffling block size $b$. A larger $b$ enhances effectiveness but compromises fidelity. For low-resolution datasets, $b=2$ provides a good trade-off between effectiveness and fidelity. However, for ImageNet-10, it demonstrates poor performance on both effectiveness and fidelity. With $b=8$, the prediction accuracy of the unauthorized images is approximately 50%, while the accuracy of authorized images drops by over 5%.
DeepIPR [19], SSAT [21], and DSN [26] can keep the accuracy drop to lower than 1.5% on authorized images and sufficiently degrade the performance on unauthorized images (<40%). However, these three methods have their own security vulnerability. The passport layers of DeepIPR can be easily localized since they require additional inputs. As demonstrated in [38], the attacker can replace these passport layers with custom normalization layers and re-train them to achieve high accuracy on benign images with a forged passport. SSAT assigns ground truth labels to poisoned training samples and wrong labels to clean training samples. The ground-truth label $y$ of clean input image is mapped to the wrong label $y^{\prime}=y+1$. This fixed bijective mapping can be easily reverse-engineered by observing the model’s outputs to a small set of labeled test images. Any other deterministic bijective mapping is as vulnerable. The DSN method directly pastes a fixed binary pattern on the test images as the secret key, which can be easily spotted by human eyes.
IDEA achieves good effectiveness and fidelity on all datasets. Layer substitution attack is not able to break IDEA as we make no modification to the model architecture. Query-based reverse engineering is also infeasible as the outputs of unauthorized images are randomized. Since the perturbations introduced by SteganoGAN encoder are imperceivable and sample-specific, it is impossible for the attacker to recognize and copy the secret key.
Although the elementary-distilled model can achieve similar classification accuracy as the protected model, their differences will be more closely scrutinized in Sec. VI.
@@ VI A Closer Look at Latent Representations
In this section, the hidden layer outputs of the unprotected, MoE, elementary-distilled and protected models are examined from three perspectives: MI estimation, attention map visualization, and representation visualization.
@@ VI-A MI Estimation
Fig. 7 visualizes the estimated MI between the latent representations collected from paired source (authorized) and target (unauthorized) test samples, with the first row showing the MI between authorized and benign images, and the second row showing the MI between authorized and noise images. Layers 1 to 4 are the selected layers for MI minimization and multi-layer attention and contrastive representation knowledge distillation, progressing from the shallow to the deep layer.
For all selected layers, we observe that the estimated MI on the unprotected model is consistently high. This is expected, as this model has not been trained to recognize steganographic perturbations and thus produces similar representations for all types of images. In contrast, the MI between the source and target images estimated on the MoE model is significantly lower than that on the unprotected model, demonstrating that the training objectives in Eqs. (12) and (13) effectively increase the expected risk of the two fake experts. The protected model also shows similarly low MI values, indicating that the protected model faithfully emulates the MoE model’s behavior across all selected layers. Consequently, if the input images are not embedded with the correct key, the adversary can only obtain meaningless intermediate layer representations that are comparable to those output by the fake experts. However, the MI estimated on the elementary-distilled model is higher than that of the MoE and protected models, and in some cases even higher than that of the unprotected model. This indicates that using only KL-divergence constraints on the output layer during knowledge distillation is insufficient to protect the model functionalities. Specifically, the shallow layers of the elementary-distilled model exhibit good generalizability on the target domains, making it easier to break the active authorization protection through fine-tuning, as these well-performed shallow layers require almost no adjustment.
Figure 8: Visualization of attention maps generated by the original unprotected, MoE, protected and elementary-distilled models, respectively. The test images are randomly selected from ImageNet-10.
@@ VI-B Attention Map Visualization
Fig. 8 shows the attention maps generated at the output of selected layers by feeding paired benign/authorized/noise test images into the four models. It shows that the unprotected model outputs similar attention maps for all three domains of test images at all selected layers. Conversely, the MoE model produces similar attention maps to the unprotected model only on input of authorized images. On input of benign or noise images, the MoE model exhibits attention behaviors akin to an untrained model to these inputs. Moreover, the protected model perfectly imitates the intermediate layer behaviors of the MoE model, as it outputs almost identical attention maps as those of the MoE model for the same domains of test images at all the selected layers. However, the elementary-distilled model exhibits similar attention maps for all types of images at the first three selected layers, and these attention maps are similar to those of the unprotected model at the same layers.
Figure 9: t-SNE visualization of the latent representations of the original unprotected, elementary-distilled, MoE and protected models obtained from 1,500 ImageNet-10 test images. The latent representations of the benign/authorized/noise test images are shown as blue squares, green dots, and red triangles, respectively.
@@
We randomly selected 500 benign images from the ImageNet-10 dataset and generated their authorized and noise versions to create a mixed test set with 1,500 images. These test images were separately fed into the four models, and their outputs at the four selected layers were collected. The latent representations of the three types of test images are visualized using the t-SNE [46] in Fig. 9. For the unprotected model, the points corresponding to the paired benign/authorized/noise images almost completely overlap across all selected layers. This observation further illustrates that the source and target images have high estimated MI at the hidden layers. The elementary-distilled model has similar behaviors in layers 1 to 3, but the representations in layer 4 form three distinct clusters. This result, again, reveals that the elementary-distilled model has good generalizability at the shallow layers, and the representations of the three types of images are indistinguishable until the deep layers. On the contrary, for the MoE and protected models, the points corresponding to benign/authorized/noise images are well clustered across all selected layers. This indicates that the MI minimization has successfully made the two fake experts behave differently from the real expert, and the protected model effectively mimics the behavior of the MoE model at not only the output layer but also the selected hidden layers.
@@ VII Security Analysis
We further investigate the robustness of IDEA against both model transformation and reverse engineering attacks. Model transformation attacks aim to remove the active authorization protection by directly modifying the pirated model through fine-tuning, model pruning, and transfer-learning. reverse engineering attacks aim to reverse engineer the encoding process without modifying the pirated model.
(a) CIFAR-10
(b) Caltech-101
(c) GTSRB
(d) ImageNet-10
Figure 10: Performance of fine-tuned models on benign test datasets. The green solid lines indicate the baseline performance of unprotected models.
VII-A Model Transformation Attacks
VII-A1 Fine-tuning
An attacker may fine-tune a stolen model to bypass active control. It is assumed that the attacker has access to only limited training data compared to the owner; otherwise, the attacker could simply train their own model from scratch.
To evaluate the robustness of IDEA against such an attack, both protected and elementary-distilled models were fine-tuned for 100 epochs using 5%, 10%, 15%, and 20% samples of the benign training dataset. Fig. 10 illustrates the accuracy of the fine-tuned models on the benign test dataset. The baseline performance of the unprotected models is represented by the green solid line. For fine-tuned protected models, the performance improves as the ratio of available samples increases. However, even with 20% training samples, the inference accuracy on benign data remains significantly lower than the baseline performance. Additionally, models fine-tuned from the elementary-distilled version exhibit higher accuracy than those fine-tuned from the protected model. This observation aligns with the analysis in Sec. VI that the shallow layers of the elementary-distilled model are more vulnerable to exploitation during fine-tuning attacks. To mitigate this, multi-layer attention and contrastive representation losses are necessary to prevent the shallow layers of the distilled model from leaking the real expert’s functionality.
TABLE VI: Accuracy (%) of models trained using transfer-learning from unprotected and protected models. The original training datasets are listed in the first row and the private dataset to be adapted in the first column. ImageNet-10$\dagger$ is another subset of ImageNet consisting of 10 classes that do not overlap with ImageNet-10.
Dataset CIFAR-10 Caltech-101 GTSRB ImageNet-10
unprotected protected unprotected protected unprotected protected unprotected protected
CIFAR-10 - - 37.35 26.05 50.85 22.80 - -
Caltech-101 68.17 54.62 - - 63.06 43.95 - -
GTSRB 46.78 35.43 31.99 20.00 - - - -
ImageNet-10 - - - - - - 50.34 40.36
@@ VII-A1 Fine-tuning
VII-A2 Model pruning
Weight pruning (WP) and filter pruning (FP), are commonly used to reduce model size by removing parameters that contribute insignificantly to inference performance [59]. WP globally prunes a fraction (ranging from 5% to 95%, with a step size of 5%) of the smallest magnitude weights. FP iteratively prunes filters in a specific convolutional layer, starting from the smallest magnitude to the largest. Following the configuration in [60], we select the last convolutional layer of ResNet-18 as the target layer for FP implementation.
@@ VII-A2 Model pruning
VII-A3 Transfer-learning
This attack aims to utilize the well-trained feature extractor to adapt the pirated model to a different task using a private dataset. The common transfer-learning paradigm freezes all layers except the last fully connected layer, which is randomly initialized based on the new task. For the four low-resolution datasets, we perform transfer learning on both unprotected and protected models using other low-resolution datasets. For the high-resolution ImageNet-10, we select another 10 classes from the entire ImageNet dataset to form a non-overlapping subset ImageNet-10$\dagger$ for transfer-learning. The results in Tabel VI show that models that are transfer-learnt from protected models exhibit worse performance than from unprotected models as protected models output meaningless features by imitating fake experts.
TABLE VII: Accuracy (%) of pirated models on reverse-engineered images generated based on Assumptions 1 and 2.
@@ VII-A3 Transfer-learning
VII-B Reverse engineering Attacks
@@
Assumption 1: The attacker only has a small portion of the benign images to train a surrogate generator to produce encoded images for the pirated model.
@@
Assumption 2: The attacker has not only a small portion of benign images but also their authorized versions embedded with the correct key by eavesdropping on legitimate users. This scenario enables the attacker to train the generator more efficiently by simultaneously minimizing the MSE loss between the generated and authorized images.
@@
We simulate these attacks by training a generator with different numbers of benign images or benign-authorized pairs, ranging from 500 to 5,000. The generator has a similar architecture to the SteganoGAN encoder but without the key input port. Both its input and output are 3-channel color images of the same size. Table VII reports the accuracy of the pirated model on test images created by the generator. Under Assumption 1, most reverse engineering attacks result in low prediction accuracy close to random guessing. The best attack performance achieved on Caltech-101 with 5,000 available benign images is merely 22.13%. As expected, the attack performance under Assumption 2 is much better in most cases, with an average improvement of 9.55%. However, even with 5,000 available benign-authorized pairs, which are difficult for the attacker to collect in reality, the best attack performance of 56.84% achievable on CIFAR-10 is still far lower than the baseline. Therefore, IDEA is resilient against even the strongest assumption of reverse engineering attacks.
VIII Limitations
Security comes at a cost. Notwithstanding the robust proactive protection for DNNs offered by IDEA, there are inevitable trade-offs.
One premium to pay for the active control is the computational cost of generating protected models. Each instance requires training three experts and performing knowledge distillation to integrate their functionalities into a unified network. This limitation is similarly observed in existing active IP protection mechanisms [25, 21, 18, 26, 19, 22, 23] designed for multiple users. Nonetheless, this process is faster than training a model from scratch, and the overhead remains manageable. For the distributed revenue business model, this cost adds only incrementally to the bracket price to purchase a customer-premise model for unlimited private use. Another aspect is the number of queries required for ownership verification, which scales linearly with the number of users. Even with a small query dataset, such as 100 samples, encoding these with all recorded keys introduces additional computational steps. However, this is a one-time process for verifying ownership and identifying dishonest users. The overhead is trivial relative to the strong protection it offers. Lastly, encoding test images through the SteganoGAN encoder introduces a slight latency during offline usage.
IX Conclusion